# Quenched coalescent for diploid population models with selfing and overlapping generations

### v1

@@ 4 Quenched convergence of the ancestral process
Briefly, let $\lambda$ and $Q$ be given by Assumption 3.4 and Remark 3.8 respectively. Then $\mathcal{L}^{N,n}$ converges in distribution to a random element $\mathcal{L}_{Q,\lambda}^{n}\in\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ described as follows: Consider a random graph $G_{Q,\lambda}^{n}=\big(G_{Q,\lambda}^{n}(t)\big)_{t\in\mathbb{R}_{+}}$, to be called a $Q$-$\lambda$ graph starting with $n$ nodes at $t=0$, in which each node in the graph splits into two nodes independently at rate $\lambda$, and where any subset of nodes coagulate into one node according to the rate matrix $Q$ (to be made precise in Definition 7.1); see Figure 4 for an illustration. Given $G_{Q,\lambda}^{n}$, we consider a system of coalescing random walks on it, where each random walker chooses one of the two directions to go when it reaches a point of fragmentation. Then $\mathcal{L}_{Q,\lambda}^{n}$ is the conditional law of this system of coalescing random walks given $G_{Q,\lambda}^{n}$; we make this precise in Definition 7.3.
@@ 4.1 Random walks on exchangeable fragmentation-coalescence processes
An exchangeable fragmentation-coalescence (EFC) process $\Pi$ is a $\mathcal{E}_{\infty}$-valued exchangeable Markov process so that the restriction to $\mathcal{E}_{n}$, denoted by $\Pi^{n}$, is a cádlág finite state Markov chain which can only evolve by fragmentation of one block or by coalescence.
By Ber (04)[p. 781] $\Pi$ is fully characterized by $\Pi(0)$, finite measures $\nu_{Disl}$ and $\nu_{Coag}$ on $\Delta$, and non-negative constants $c_{k},c_{e}$. In this paper we will always consider the case $\nu_{Disl}=0$. $c_{k}$ and $\nu_{Coag}$ form the coalescing part of the EFC. Specifically, for any finite subset $l$ of particles in the EFC, they coalesce as if they were an $l$-$\Xi$-coalescent where $\Xi=c_{k}\delta_{\mathbf{0}}+\nu_{Coag}$. That is, the non-Kingman coalescences are governed by $\nu_{Coag}$ and $c_{k}$ is the rate at which any pair of particles coalesce. $c_{e}$ is simply the rate at which each particle in the EFC fragments into two particles.
Example 4.2 (Ancestral Recombination Graph).
Suppose $\Pi=\left(\Pi(t)\right)_{t\in\mathbb{R}_{+}}$ is an EFC with characteristics $c_{k}=2$, $\nu_{Coag}=0$, $c_{e}=\lambda,$ and $\nu_{Disl}=0$. Then for any $n$ particles, each of the $\binom{n}{2}$ pairs of particles coalesce independently with rate $2$ and each of the $n$ particles fragment independently at rate $\lambda$, which is preceisely the structure of an ancestral selection or ancestral recombination graph Gri (91); GM (97); KN (97). This subgraph of $\Pi$ was proposed in (NWF, 25, Section 4.1) as the scaling limit of a diploid Moran model with selfing. This is established in Section 5.1.
@@
Suppose that $\psi$ is a fixed constant in $[0,1]$. By a $(\psi,\lambda,\rho)$-EFC $\Pi(\psi,\lambda,\rho)$ we mean an EFC where $c_{k}=2$, $c_{e}=\lambda$, and
$\nu_{Coag}=\frac{\psi^{2}}{2}\delta_{\left(\frac{\psi}{2},\frac{\psi}{2},0,0,\ldots\right)}.$
This EFC has implicitly appeared in BBE (13) when $\psi$ is a fixed constant.
@@
Suppose that $\psi$ is random and Beta(2r,r)(2-r,r) distributed for some $r\in[0,1]$. Let $\nu_{Coag}$ denote the measure on $\Delta\setminus\{\mathbf{0}\}$ defined by
$\int_{0}^{1}\frac{z^{2}}{2}\delta_{\left(\frac{z}{2},\frac{z}{2},0,0,\ldots\right)}\,{\rm Beta}(2-r,r)(dz),,$
@@
There is a three-parameter family of EFCs $\Pi(r,\rho,\lambda)$ with characteristics $c_{k}=2$, $\nu_{\rm Coag}=\rho\Xi^{\prime}(\psi)$, $c_{e}=\lambda$, and $\nu_{\rm Frag}=0$ that arise as the scaling limit of models where a highly reproductive couple gives birth to $\psi\simBeta$(2r,r)(2-r,r) of the population, as in BBE (13). This is established in Section 5.4.
@@
The ancestral recombination graph of BBE (13) is in fact the subgraph of the $\Pi$ of Example 4.3 given by tracing all trajectories of the first $n$ blocks of $\Pi$ backwards in time.
@@
Figure 4: Here we see on a fixed realization of an EFC $\Pi$ with $c_{k}=1$ two realizations of coalescing random walks on $\Pi$. These random walks follow each coalescence in the EFC and choose between each of the possible edges ahead of them at a fragmentation with equal likelihood.
@@
We provide, in the main result of this paper contained in Section 4.2 criteria by which $\mathcal{L}_{\Pi}^{n}$ is the quenched scaling limit of the law of the coalescent conditional on the pedigree.
@@ 4.2 Main result
We are now ready to state our main result in this paper. Recall that $\mathcal{L}^{N,n}:=\mathbb{P}\left(\bar{\chi}^{N,n}\in\cdot\mid\mathcal{A}_{N}\right)$ was defined in (4.1), and $\mathcal{L}_{\Pi}^{n}$ was defined in Definition 4.6.
@@
Suppose Assumptions 3.3, 3.6 and 3.4 hold. Then there is a unique measure $\Xi$ on $\Delta$, as explained in Remark 3.8 such that, for $\Pi$ an EFC with characteristics $c_{k}=\Xi(\mathbf{0})$, $\nu_{Coag}=\nu_{\Xi-c_{k}\delta_{\mathbf{0}}}$, $c_{e}=\lambda$, and $\nu_{Disl}=0$, it holds that $\mathcal{L}^{N,n}$ converges in distribution to $\mathcal{L}_{\Pi}^{n}$ in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ as $N\to\infty$.
Theorem 4.7 says that when the outcrossing timescale and the coalescence timescale are comparable, the structure of the $n$ sample lineages, conditional on the pedigree, converges to that of coalescing random walks on a realization of an EFC process.
@@
Note that the form of weak convergence in distribution is stronger than that described by the Meyer-Zheng topology of MZ (84). In particular, it follows from Theorem 4.7 that the random occupation measure in $\mathcal{M}_{\rm loc}\left(\mathbb{R}_{+}\times\mathcal{E}_{n}\right)$ defined by $dt\otimes\delta_{\bar{\chi}^{N,n}(t)}$ converges vaguely in distribution with respect to the measure $\mathbb{P}\left(\cdot\,\mid\,\mathcal{A}_{N}\right)$ to the random measure $dt\otimes\delta_{\chi^{n}(t)}$ governed by $\mathbb{P}\left(\cdot\,\mid\,\Pi\right)$.
@@
[Relaxing assumption on sampling]In our main result, we assumed that the $n$ sampled gene copies are from $n$ distinct individuals, one gene copy from each individual. This assumption can be relaxed and the corresponding result can be obtained by our method, and we briefly describe the corresponding limiting object here. Suppose the $n$ sampled gene copies consist of $2m$ gene copies sampled in pairs within individuals and n2mn-2m gene copies sampled each from distinct individuals. Then, since $\alpha_{N}\to 1$, the two gene copies within each of the $m$ pairs will instantaneously coalesce with probability 1. Both the annealed limit and the quenched limit (as $N\to\infty$) will be the same as before, but starts with nmn-m gene copies rather than $n$. The annealed limit in this case (for fixed selfing probability) was obtained in ND (97). For the proof, one may specify what it means to take supremum in an empty set in 3.4 and show that the transient behavior of the discrete-time processes in $\mathcal{S}_{n}$ converge to that of the limiting process.
@@ 5 Applications
In this section, we discuss some important applications of our main result that motivate this study.
In Section 5.1 we establish that the coalescing random walk on the classical ancestral recombination graph (ARG), first introduced for a diploid Moran model with selfing in (NWF, 25, Section 4.1), arises as the limiting conditional coalescent for the corresponding diploid Wright-Fisher model, and to perturbations of both of these models as well.
In Section 5.2, we prove that the form of convergence for $\mathcal{L}^{N,n}$ implies the convergence of important tree statistics in population genetics, in particular the branch lengths associated with the site-frequency spectrum, or SFS. We also illustrate how the SFS depends on the pedigree when the selfing probability is close to one.
In Section 5.2, we establish that some important tree statistics in population genetics converge. Of particular importance in population genetics are certain classes of integral functionals, namely the internal and external branch lengths of the coalescent process. These statistics form the basis of inference methods in population genetics, and so we need to know that the form of convergence for $\mathcal{L}^{N,n}$ implies the convergence of these statistics. The adequate notion of convergence necessary is weak convergence in distribution of random measures, which we make precise in Section 6. The standard reference is (Kal, 17, Chapter 4).
@@
5.1 Robustness of coalescing walks on the ancestral recombination graph
@@
Suppose that Assumptions 3.3, 3.4 and 5.1 hold. Then Assumption 3.6 holds and the sequence of random measures $\mathbb{P}_{\xi_{0}^{n}}\left(\bar{\chi}^{N,n}\in\cdot\,\mid\,\mathcal{A}_{N}\right)$ converge weakly in distribution to $\mathcal{L}_{\Pi}^{n}$ where $\Pi$ is the ancestral recombination graph described in Example 4.2.
@@
By Lemma Appendix C.2, $p_{\xi\xi}^{N,n}=1-2\binom{|\xi|}{2}c_{N}+o(c_{N})$. Therefore the limiting exit rate from the state $\xi$ is $2\binom{\xi}{2}$ when we take the $c_{N}$ time-rescaling with $c_{N}\to 0$. Further, Lemma Appendix C.2 demonstrates that $\mathit{p}_{\xi\eta}^{N,n}\in o(c_{N})$ for any $\eta$ that cannot be obtained from $\xi$ via a single binary merger, i.e. for which $\xi\not\prec\eta$. For such $\xi,\eta$ we have that the limiting exit rate from $\xi$ to $\eta$ is 0 with the $c_{N}$ time-rescaling with $c_{N}\to 0$. Finally, there are $\binom{|\xi|}{2}$ partitions $\eta$ for which $\xi\prec\eta$. By exchangeability, these are equally likely and so the limiting transition rate from $\xi$ to $\eta$ is $2$ when $\xi\prec\eta$.
@@
$Q_{n}$ is the transition-rate matrix of a Kingman $n$-coalescent with time rescaled by a factor of $2$. Hence Assumption 3.6 is satisfied and $(Q_{n})_{n\in\mathbb{N}}$ determines a measure $\Xi=2\delta_{\mathbf{0}}$ on $\Delta$ in the sense of Remark 3.8. As Assumptions 3.3 and 3.4 hold, and we have shown that $Q_{N,n}$ converges to the generator of a time-rescaled Kingman $n$-coalescent, the claim follows directly from Theorem 4.7. ∎

### v2

@@ 4 Quenched convergence of the ancestral process
Briefly, let $\lambda$ and $Q$ be given by Assumption 3.5 and Remark 3.10 respectively. Then $\mathcal{L}^{N,n}$ converges to a random element $\mathcal{L}_{Q,\lambda}^{n}\in\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ described as follows: Consider a random graph $G_{Q,\lambda}^{n}=\big(G_{Q,\lambda}^{n}(t)\big)_{t\in\mathbb{R}_{+}}$, to be called a $Q$-$\lambda$ graph starting with $n$ nodes at $t=0$, in which each node in the graph splits into two nodes independently at rate $\lambda$, and where any subset of nodes coagulate into one node according to the rate matrix $Q$ (to be made precise in Definition 7.1); see Figure 4 for an illustration. Given $G_{Q,\lambda}^{n}$, we consider a system of coalescing random walks on it, where each random walker chooses one of the two directions to go when it reaches a point of fragmentation. Then $\mathcal{L}_{Q,\lambda}^{n}$ is the conditional law of this system of coalescing random walks given $G_{Q,\lambda}^{n}$; we make this precise in Definition 7.3.
@@ 4.1 Random walks on exchangeable fragmentation-coalescence processes
An exchangeable fragmentation-coalescence (EFC) process $\Pi=(\Pi(t))_{t\in\mathbb{R}_{+}}$ is a $\mathcal{E}_{\infty}$-valued exchangeable Markov process so that the restriction to $\mathcal{E}_{n}$, denoted by $\Pi^{n}$, is a cádlág finite state Markov chain which can only evolve by fragmentation of one block or by coalescence.
By (Ber, 04, p. 781) $\Pi$ is fully characterized by the initial state $\Pi(0)$, finite measures $\nu_{Disl}$ and $\nu_{Coag}$ on $\Delta$ where $\nu_{Coag}(\mathbf{0})=0$, and non-negative constants $c_{k}$ and $c_{e}$. In this paper we will always consider the case $\nu_{Disl}=0$. The constant $c_{e}$ is the rate at which each particle in the EFC fragments into two particles, while $c_{k}$ and $\nu_{Coag}$ form the coalescing part of the EFC. Specifically, for any finite subset $l$ of particles in the EFC, they coalesce as if they were an $l$-$\Xi$-coalescent where
$\Xi=c_{k}\delta_{\mathbf{0}}+\nu_{Coag}.$ (4.2)
That is, the non-Kingman coalescences are governed by $\nu_{Coag}$ and $c_{k}$ is the rate at which any pair of particles coalesce independently.
Example 4.2 (Ancestral Graphs).
Suppose $\Pi=\left(\Pi(t)\right)_{t\in\mathbb{R}_{+}}$ is an EFC with characteristics $c_{k}=2$, $\nu_{Coag}=0$, $c_{e}=\lambda,$ and $\nu_{Disl}=0$. Then for any $n$ particles, each of the $\binom{n}{2}$ pairs of particles coalesce independently with rate $2$ and each of the $n$ particles fragment independently at rate $\lambda$, which is precisely the structure of an ancestral recombination graph or an ancestral selection graph Gri (91); GM (97); KN (97), aside from the factor-of-$2$ difference in time scale noted in Remark 3.12. This subgraph of $\Pi$ was conjectured in (NWF, 25, Section 4.1) as the scaling limit of a diploid Moran model with limited outcrossing. This is established in Section 5.1 in this paper.
@@
Suppose that $\psi\in[0,1]$ and $\rho\in(0,\infty)$ are fixed constants. By a $(\psi,\lambda,\rho)$-EFC, denoted by $\Pi(\psi,\lambda,\rho)$, we mean an EFC where $c_{k}=2\frac{1}{1+\rho\frac{\psi^{2}}{2}}$, $c_{e}=\frac{\lambda}{1+\rho\frac{\psi^{2}}{2}}$, and
$\nu_{Coag}(dx)=2\frac{\rho\frac{\psi^{2}}{2}}{1+\rho\frac{\psi^{2}}{2}}\delta_{\left(\frac{\psi}{2},\frac{\psi}{2},0,0,\ldots\right)}(dx).$
This EFC arises from the model in BBE (13) when $\psi\in[0,1]$ is a fixed constant, as we proved in Corollary 5.7.
@@
Consider the EFC with characteristics $c_{k}=2\frac{1}{1+\frac{1}{12}\frac{\Gamma(4-r)}{\Gamma(2-r)}}$, $c_{e}=\lambda\frac{1}{1+\frac{1}{12}\frac{\Gamma(4-r)}{\Gamma(2-r)}}$, $\nu_{Disl}=0$ and
$\nu_{Coag}=2\frac{\frac{\rho}{12}\frac{\Gamma(4-r)}{\Gamma(2-r)}}{1+\frac{\rho}{12}\frac{\Gamma(4-r)}{\Gamma(2-r)}}\rho\int_{0}^{1}\frac{z^{2}}{2}\delta_{\left(\frac{z}{2},\frac{z}{2},0,0,\ldots\right)}\,{\rm Beta}(2-r,r)(dz),,$
@@
This EFC arise when $\psi$ in Example 4.3 is random and Beta(2r,r)(2-r,r) distributed for some $r\in[0,1]$.
@@
For $\Pi$ the EFC of Example 4.3, with an appropriate time-rescaling, the subgraph given by tracing the possible trajectories of coalescing random walks starting at the first $n$ blocks of $\Pi$ backwards in time is equal in distribution to the ancestral recombination graph of BBE (13). In particular, the gene genealogies given by coalescing random walks on $\Pi$ are the same as for the gene genealogy of a single locus in BBE (13).
@@
Figure 4: Two realizations of coalescing random walks on a single, fixed realization of an EFC $\Pi$ with $c_{k}=1$. These random walks follow each coalescence in the EFC and choose between each of the possible edges ahead of them at a fragmentation with equal probability.
@@
We provide, in the main result of this paper contained in Section 4.2, criteria by which $\mathcal{L}_{\Pi}^{n}$ is the quenched scaling limit of the law of the coalescent conditional on the pedigree.
@@ 4.2 Main result
We are now ready to state our main result. Recall that $\mathcal{L}^{N,n}:=\mathbb{P}\left(\bar{\chi}^{N,n}\in\cdot\mid\mathcal{A}_{N}\right)\in\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ was defined in (4.1), and $\mathcal{L}_{\Pi}^{n}$ was defined in Definition 4.6. Throughout this paper, we equip the space $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ with the weak topology. Under this topology, convergence in distribution (denoted by $\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}$}\, in this paper) of a sequence $(\mu_{N})_{N\in\mathbb{N}}$ of random elements in $\mathcal{M}_{1}\left(\mathcal{S}\right)$ is also called “weak convergence in distribution” in (Kal, 17, Chapter 4).
@@
Suppose Assumptions 3.4, 3.7 and 3.5 hold. Then there is a positive finite measure $\Xi$ on $\Delta$ (as explained in Remark 3.10) such that, for $\Pi$ an EFC with characteristics $c_{k}=\Xi(\mathbf{0})$, $\nu_{Coag}=\nu_{\Xi-c_{k}\delta_{\mathbf{0}}}$, $c_{e}=\lambda$, and $\nu_{Disl}=0$, it holds that $\mathcal{L}^{N,n}$ converges weakly in distribution to $\mathcal{L}_{\Pi}^{n}$ in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ as $N\to\infty$. That is,
$\mathcal{L}^{N,n}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mathcal{L}_{\Pi}^{n}\quad\text{in}\quad\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right).$
The proof of Theorem 4.7 is contained in Section 7. The proof proceeds by analyzing the possible trajectories of sample lineages in the pedigree backwards in time. On this graph of trajectories the sample lineages approximately perform a family of coalescing simple random walks. We show that the trace of these trajectories converge to a subgraph of a suitable EFC, and by a continuity argument show that $\mathcal{L}^{N,n}$ converges weakly in law to the random law governing the coalescing random walks on that EFC.
@@
Compared to Theorem 3.11 which says that the (unconditional) law of $\bar{\chi}^{N,n}$ converges to a deterministic element (the law of a $\Xi$-coalescent) in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$, Theorem 4.7 says that the conditional law $\mathcal{L}^{N,n}:=\mathbb{P}\left(\bar{\chi}^{N,n}\in\cdot\mid\mathcal{A}_{N}\right)$ converges to a random element in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ and the randomness in the limit is entirely captured by the EFC $\Pi$ for all sample size $n\geq 2$.
@@
Note that the form of weak convergence, made precise in Section 6, is stronger than that described by the Meyer-Zheng topology of MZ (84). In particular, it follows from Theorem 4.7 that the random occupation measure in $\mathcal{M}_{\rm loc}\left(\mathbb{R}_{+}\times\mathcal{E}_{n}\right)$ defined by $dt\otimes\delta_{\bar{\chi}^{N,n}(t)}$ converges vaguely in distribution with respect to the measure $\mathbb{P}\left(\cdot\,\mid\,\mathcal{A}_{N}\right)$ to the random measure $dt\otimes\delta_{\chi^{n}(t)}$ governed by $\mathbb{P}\left(\cdot\,\mid\,\Pi\right)$.
Remark 4.10 (Relaxing assumption on sampling).
In our main result, we assumed that the $n$ sampled gene copies are from $n$ distinct individuals, one gene copy from each individual. This assumption can be relaxed and the corresponding result can be obtained by our method, and we briefly describe the corresponding limiting object here. Suppose the $n$ sampled gene copies consist of $2m$ gene copies sampled in pairs within individuals and n2mn-2m gene copies sampled each from distinct individuals. Then, since $\alpha_{N}\to 1$, the two gene copies within each of the $m$ pairs will instantaneously coalesce with probability 1. Both the annealed limit and the quenched limit (as $N\to\infty$) will be the same as before, but starts with nmn-m gene copies rather than $n$. The annealed limit in this case (for fixed selfing probability) was obtained in ND (97). For the proof, one may specify what it means to take supremum in an empty set in 3.4 and show that the transient behavior of the discrete-time processes in $\mathcal{S}_{n}$ converge to that of the limiting process.
@@ 5 Applications
In this section, we discuss some applications of our main result that motivate this study.
In Section 5.1 we establish that the system of coalescing random walks on the ancestral graph with binary mergers (like the ancestral recombination graph) arises as the limiting conditional coalescent for a class of models including both the Wright-Fisher model and the Moran model, showing robustness of such system. This result extends that of NWF (25) from sample of size 2 to general size $n\geq 2$.
In Section 5.2, we prove that our main theorem implies the convergence of important tree statistics in population genetics as $N\to\infty$. We consider the internal and external branch lengths of the coalescent tree associated with the site-frequency spectrum (SFS). We also illustrate, through simulation of the robust model in Section 5.1, how the SFS depends on the pedigree when the selfing probability is close to one.
@@
5.1 Robustness of coalescing walks on the ancestral graph with binary mergers
@@
Suppose that Assumptions 3.4, 3.5 and 5.1 hold. Then Assumption 3.7 holds and the sequence of random measures $\mathbb{P}_{\xi_{0}^{n}}\left(\bar{\chi}^{N,n}\in\cdot\,\mid\,\mathcal{A}_{N}\right)$ converge weakly in law to $\mathcal{L}_{\Pi}^{n}$ where $\Pi$ is the ancestral recombination graph described in Example 4.2.
@@
By Lemma Appendix C.2, $h_{\xi\xi}^{N,n}=1-2\binom{|\xi|}{2}c_{N}+o(c_{N})$. Therefore, the limiting exit rate of the process $\bar{\chi}^{N,n}$ from the state $\xi\in\mathcal{E}_{n}$ is $2\binom{\xi}{2}$ as $N\to\infty$, by Assumptions 3.4. Further, Lemma Appendix C.2 demonstrates that $h_{\xi\eta}^{N,n}\in o(c_{N})$ for any $\eta\in\mathcal{E}_{n}$ that cannot be obtained from $\xi$ via a single binary merger, i.e. for which $\xi\not\prec\eta$. For such $\xi,\eta$ we have that the limiting exit rate from $\xi$ to $\eta$ is 0 by Assumptions 3.4. Finally, there are $\binom{|\xi|}{2}$ partitions $\eta$ for which $\xi\prec\eta$. By exchangeability, these are equally likely and so the limiting transition rate from $\xi$ to $\eta$ is $1$ when $\xi\prec\eta$.
@@
$Q_{n}$ is the transition-rate matrix of a Kingman $n$-coalescent with time rescaled by a factor of $2$. Hence Assumption 3.7 is satisfied and $(Q_{n})_{n\in\mathbb{N}}$ determines a measure $\Xi=2\delta_{\mathbf{0}}$ on $\Delta$ in the sense of Remark 3.10. As Assumptions 3.4 and 3.5 hold, and we have shown that $Q_{N,n}$ converges to the generator of a time-rescaled Kingman $n$-coalescent, the claim follows directly from Theorem 4.7. ∎